# SPECIAL RELATIVITY

### v9

@@ 15.2 Paraxial Approximation for the Radiation Field
The general method to derive the frequency spectrum is to transform the electric field from the time domain to the frequency domain by the use of Fourier transform. First, let us rewrite Eq. (LABEL:revwied) as follows
@@
Eq. (LABEL:revwied) and Eq. (LABEL:rrevwied) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral (see Appendix I).
We call $z_{0}$ the observation distance along the optical axis of the system, while $(x_{0},y_{0})$ fixes the transverse position of the observer. Using the complex notation, in this and in the following sections we assume, in agreement with Eq. (28), that the temporal dependence of fields with a certain frequency is of the form:
@@
In the following we will always assume that the ultra-relativistic approximation is satisfied, which is the case for SR setups. As a consequence, the paraxial approximation applies too. The paraxial approximation implies a slowly varying envelope of the field with respect to the wavelength. It is therefore convenient to introduce the slowly varying envelope of the transverse field components as
@@
We will now replace all vectors by their components to obtain directional dependency of the synchrotron radiation. The emission angle $\theta=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ is taking with respect to the $z$-axis. Here $\theta_{x}=x_{0}/z_{0}$ is the observation angle projected onto the $x-z$ plane, $\theta_{y}=y_{0}/z_{0}$ is the observation angle projected onto the $y-z$ plane. The components of the unit vector $\vec{n}$ can be approximated by $n_{z}=1-\theta_{x}^{2}/2-\theta_{y}^{2}/2$, $n_{x}=\theta_{x}$, $n_{y}=\theta_{y}$, so $\vec{n}\cdot(\vec{r}_{0}-\vec{r})=(z_{0}-z)(1-\theta_{x}^{2}/2-\theta_{y}^{2}/2)-x\theta_{x}-y\theta_{y}$. We consider the motion in a static magnetic field. According to the conventional particle tracking the magnitude of the velocity is constant and is equal $v=ds/dt=\mathrm{const.}$, where $s(z)$ is the longitudinal coordinate along the path. The transverse components of the envelope of the field in Eq. (LABEL:rrevwied) in the far zone and in paraxial approximation finally becomes
@@ 15.3 Undulator Radiation
Traditionally, all courses in synchrotron radiation theory have begun in the same way, retracing the path followed in the historical development of the subject. One first learns about synchrotron radiation from bending magnet. We will begin in this chapter by dealing with the ”advanced” subject of the synchrotron radiation theory. For undulator setup covariant calculations of the radiation process is quite simple.
To generate specific synchrotron radiation characteristics, radiation is often produced from special insertion devices called undulators. The resonance approximation, that can always be applied in the case of undulator radiation setups, yields simplifications of the theory. This approximation does not replace the paraxial one, but it is used together with it. It takes advantage of another parameter that is usually large, i.e. number of undulator periods $N_{w}\gg 1$. The frequency emitted by a particle going through an undulator can be obtained by considering the interference between the parts of the radiation created at successive periods (see Fig. 65). The frequency of this field is Doppler shifted. The shortest wavelength is observed on undulator axis. In resonance approximation, all undulator radiation at shortest wavelength is emitted within an angle much smaller than $1/\gamma$ (see Fig. 66). This automatically selects observation angles of interest. In fact, if we consider observation angles outside the diffraction angle, we obtain zero intensity with accuracy $1/N_{w}$.
We now have to ask: Why did the error in insertion device theory remain so long undetected? We answer this question in great detail bellow. In discussing this we will deal only with the radiation into the central cone (see Fig. 66). In doing so, we neglected all information about the electron trajectory. In this approximation the electron oscillation amplitude is always (independently of how large the undulator strength parameter) much smaller than the radiation diffraction size at the undulator exit, because $N_{w}\gg 1$. Thus, the undulator radiation theory gives fields very much like the (dipole-like) instantaneous theory. So, we are satisfied using the conventional (i.e. non-covariant) approach for describing the undulator radiation into the central cone, that is the practical situation of interest.
@@ 15.3.1 Conventional Theory
Eq. (38) can be used to characterize the far field from an electron moving on any path. In this section we present a simple derivation of the frequency representation of the radiated field produced by an electron in the planar undulator. The magnetic field on the undulator axis has the form
@@
$m$ being the electron mass at rest and $B_{w}$ being the maximal magnetic field of the undulator on axis.
In this case the electron path is given by
@@
Usually, it does not make sense to calculate the intensity distribution from Eq. (46) alone, without extra-terms (both interfering and not) from the other parts of the electron path. This means that one should have complete information about the electron path and calculate extra-terms to be added to Eq. (46) in order to have the total field from a given setup. Yet, we can find particular situations for which the contribution from Eq. (46) is dominant with respect to others. In this case Eq. (46), alone, has independent physical meaning.
@@
A well known expression for the angular distribution of the first harmonic field in the far zone (see Appendix II for a detailed derivation) can be obtained from Eq. (46). Such expression is axis-symmetric, and can, therefore, be presented as a function of a single observation angle $\theta$, where $\theta^{2}=\theta_{x}^{2}+\theta_{y}^{2}$. One obtains the following distribution for the slowly varying envelope of the electric field:
@@
where $\mathrm{sinc}(\cdot)\equiv\sin(\cdot)/(\cdot)$. Therefore, the field is horizontally polarized and azimuthal symmetric. Eq. (55) describes a field with spherical wavefront centered in the middle of the undulator.
@@ 15.3.2 Why did the Error in Insertion Device Theory Remain so Long Undetected?
We have seen that in all generality the expression for the undulator field in the far zone and in the ultrarelativistic (i.e. paraxial) approximation can be written as Eq. (152). Within the resonance approximation ($N_{w}\gg 1$) for the frequencies around the first harmonic it can be simplified to the well-known expression Eq. (55) where the field is horizontally polarized and azimuthal symmetric. The divergence of this radiation is much smaller compared to the angle $1/\bar{\gamma}_{z}$. The mathematical reason stems from the fact that the factor $\sin(\cdot)/(\cdot)$ represents the well-known resonance character of the undulator radiation. If we are interested in the angular width of the peak around the observation angle $\theta=0$, we can introduce an angular displacement $\Delta\theta$. Taking the first zero of the $\sin(\cdot)/(\cdot)$ function at $C=0$ we will be able to determine the natural angular width of the radiation for the first harmonic $\theta_{c}$. The cone with aperture $\theta_{c}$ is usually called central cone. It can be found that $\theta_{c}^{2}=1/(2N_{w}\bar{\gamma}_{z}^{2})\ll 1/\bar{\gamma}_{z}^{2}$.
@@
Now we would like to understand what is the characteristic transverse size of the field distribution at the exist of the undulator. The radiation from magnetic poles always interferes coherently at zero angle with respect to undulator axis. This interference is constructive within an angle of about $\sqrt{c/(\omega L_{w})}$. We can estimate the interference size at the undulator exit as about $\sqrt{cL_{w}/\omega}$. On the other hand, the electron oscillating amplitude is given by $r_{w}=c\theta_{s}/k_{w}=cK/(\gamma k_{w})$. It follows that $r_{w}^{2}/(cL_{w}/\omega)=K^{2}\omega/(L_{w}k_{w}^{2}\gamma^{2})=K^{2}/[\pi N_{w}(1+K^{2}/2)]\ll 1$, where we use the fact that $\gamma^{2}=(1+K^{2}/2)\bar{\gamma}_{z}^{2}$. This inequality holds independently of the value of $K$, because $N_{w}\gg 1$. Thus, the electron oscillating amplitude is always much smaller than the radiation diffraction size at the undulator exit.
@@
We consider the radiation associated with the first order term in the expansion of the Eq. (158) in power of small parameter $[K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq. (162)). But in doing so we miss all information about transverse electron trajectory in the phase factor Eq. (48) since the term $K\theta_{x}\omega\cos(k_{w}z^{\prime})/(\gamma k_{w}c)$ is neglected. In this approximation the electron orbit scale is always much smaller than the radiation diffraction size and Eq. (55) gives fields very much in agreement with the dipole radiation theory. So we are satisfied using the non covariant approach when considering the transverse electron motion.
@@
There are several points to be made about the above result. We have just explained that in accounting only for the radiation in the central cone, we miss all information about the transverse electron motion. To be complete we must add an analysis of the accelerated motion along the $z$-direction (i.e. along the undulator axis). We assume that the transverse velocity $v_{\perp}(z)$ is small compared to the velocity of light $c$. We consider the small expansion parameter $v_{\perp}/c$, neglecting terms of order $(v_{\perp}/c)^{3}$, but not of order $(v_{\perp}/c)^{2}$. In other words we use a second order relativistic approximation for transverse motion. We should remark that the analysis of the longitudinal motion in the ultrarelativistic approximation is much simpler than in the case of transverse motion. It is easy to see that the acceleration in the constant magnetic field yields an transverse electron velocity $v_{\perp}$ and $\Delta v_{z}=-v(v_{\perp}/c)^{2}/2$ parallel to the $z$-axis.
@@ 15.3.3 Influence of the Kick According to Conventional Theory
Eq. (55) can be generalized to the case of a particle with a given offset $\vec{l}$ and deflection angle $\vec{\eta}$ with respect to the longitudinal axis, assuming that the magnetic field in the undulator is independent of the transverse coordinate of the particle (see Appendix III). Although this can be done using Eq. (174) directly, it is sometimes possible to save time by getting the answer with some trick. For example, in the undulator case one takes advantage of the following geometrical considerations, which are in agreement with rigorous mathematical derivation. First, we consider the effect of an offset $\vec{l}$ on the transverse plane, with respect to the longitudinal axis $z$. Since the magnetic field experienced by the particle does not change, the far-zone field is simply shifted by a quantity $\vec{l}$. Eq. (55), can be immediately generalized by systematic substitution of the transverse coordinate of observation, $\vec{r}_{0}$ with $\vec{r}_{0}-\vec{l}$. This means that $\vec{\theta}=\vec{r}_{0}/z_{0}$ must be substituted by $\vec{\theta}-\vec{l}/z_{0}$, thus yielding
@@
Let us now discuss the effect of a deflection angle $\vec{\eta}$. Since the magnetic field experienced by the electron is assumed to be independent of its transverse coordinate, the path followed is still sinusoidal, but the effective undulator period is now given by $\lambda_{w}/\cos(\eta)\simeq(1+\eta^{2}/2)\lambda_{w}$. This induces a relative red shift in the resonant wavelength $\Delta\lambda/\lambda\sim\eta^{2}/2$. In practical cases of interest we may estimate $\eta\sim 1/\gamma$. Then, $\Delta\lambda/\lambda\sim 1/\gamma^{2}$ should be compared with the relative bandwidth of the resonance, that is $\Delta\lambda/\lambda\sim 1/N_{w}$, $N_{w}$ being the number of undulator periods. For example, if $\gamma>10^{3}$, the red shift due to the deflection angle can be neglected in all situations of practical relevance. As a result, the introduction of a deflection angle only amounts to a rigid rotation of the entire system. Performing such rotation we should account for the fact that the phase factor in Eq. (56) is indicative of a spherical wavefront propagating outwards from position $z=0$ and remains thus invariant under rotations. The argument in the $\mathrm{sinc}(\cdot)$ function in Eq. (56), instead, is modified because the rotation maps the point $(z_{0},0,0)$ into the point $(z_{0},-\eta_{x}z_{0},-\eta_{y}z_{0})$. As a result, after rotation, Eq. (56) transforms to
@@
Finally, in the far-zone case, we can always work in the limit for $l/z_{0}\ll 1$, that allows one to neglect the term ${\vec{l}}/{z_{0}}$ in the argument of the $\mathrm{sinc}(\cdot)$ function, as well as the quadratic term in $\omega l^{2}/(2cz_{0})$ in the phase. Thus Eq. (58) can be further simplified, giving the generalization of Eq. (55) in its final form:
@@ 15.3.4 Influence of the Kick According to Correct Coupling of Fields and Particles
According to the correct coupling of fields and particles, there is a remarkable prediction of undulator radiation theory concerning to the undulator radiation from the single electron with and without kick. Namely, when a kick is introduced, there is a red shift in the resonance wavelength of the undulator radiation in the velocity direction. To show this, let us consider the covariant treatment, which makes explicit use of Lorentz transformations.
@@
Suppose that without kick the electron moves along the constrained trajectory parallel to the undulator axis. The field which produces this electron in the far zone is given by Eq. (55). Referring back to the Eq. (59), we see that the conventional undulator radiation theory gives the following expression for radiation field after the kick
@@
This formula has nearly, but not quite the same form as Eq. (61), the difference consisting in the term $\omega\theta_{k}^{2}/(2c)$ in the argument of $\mathrm{sinc}$ function. Attention must be called to the difference in resonance frequency between the undulator radiation setup with and without the kick. Remembering the definition of the detuning parameter $C=k_{w}\Delta\omega/\omega_{r}$, we can write the red shift in resonance frequency as $\Delta\omega/\omega_{r}=-\omega_{r}\theta_{k}^{2}/(2k_{w}c)$. With this we also pointed out that the red shift can be written as $\Delta\omega/\omega_{r}=-\gamma^{2}\theta_{k}^{2}/(1+K^{2}/2)$. We now see a second order correction $\theta_{k}^{2}$ that is, however, multiplied by the large factor $\bar{\gamma}_{z}^{2}$.
We are now ready to investigate, more generally, what form the field expression takes under the introduction of a kick. Suppose that, without kick, the electron moves along the trajectory with angle $\vec{\eta}$ with respect to the undulator axis. The field produced by this electron is given by Eq. (59). We let $\vec{\theta}_{k}$ be the kick angle of the electron with respect to its initial motion. The conventional approach gives the following expression for the field after the kick
@@
Now this all leads to an interesting situation. According to the conventional theory, the resonance wavelength depends only on the observation angle with respect to the electron velocity direction. Equation (65) says that for any kick angle $\vec{\theta}_{k}$ and for any angle $\vec{\eta}$ between the undulator axis and the initial electron velocity direction, the radiation along the velocity direction has no red shift. We would like to emphasize a very important difference between conventional and covariant theory. The result of the covariant approach Eq. (67) clearly depends on the absolute value of the kick angle $\theta_{k}$ and the radiation along the velocity direction has the red shift only when the kick angle has nonzero value.
We must conclude that when we accelerate the electron in the lab frame upstream the undulator, the information about this acceleration is included into the covariant trajectory.

### v10

@@ 19.2 Paraxial Approximation for the Radiation Field
The general method to derive the frequency spectrum is to transform the electric field from the time domain to the frequency domain by the use of the Fourier transform. First, let us rewrite Eq. (LABEL:revwied) as follows
@@
Eq. (LABEL:revwied) and Eq. (LABEL:rrevwied) are equivalent but include different integrands. This is no mistake, as different integrands can lead to the same integral (see Appendix A1).
We call $z_{0}$ the observation distance along the optical axis of the system, while $(x_{0},y_{0})$ fixes the transverse position of the observer. Using the complex notation, in this and in the following sections we assume, in agreement with Eq. (LABEL:ftdef2), that the temporal dependence of fields with a certain frequency is of the form:
@@
In the following, we will always assume that the ultra-relativistic approximation is satisfied, which is the case for SR setups. As a consequence, the paraxial approximation applies too. The paraxial approximation implies a slowly varying envelope of the field with respect to the wavelength. It is therefore convenient to introduce the slowly varying envelope of the transverse field components as
@@
We will now replace all vectors by their components to obtain directional dependency of the synchrotron radiation. The emission angle $\theta=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ is taking with respect to the $z$-axis. Here $\theta_{x}=x_{0}/z_{0}$ is the observation angle projected onto the $x-z$ plane, $\theta_{y}=y_{0}/z_{0}$ is the observation angle projected onto the $y-z$ plane. The components of the unit vector $\vec{n}$ can be approximated by $n_{z}=1-\theta_{x}^{2}/2-\theta_{y}^{2}/2$, $n_{x}=\theta_{x}$, $n_{y}=\theta_{y}$, so $\vec{n}\cdot(\vec{r}_{0}-\vec{r})=(z_{0}-z)(1-\theta_{x}^{2}/2-\theta_{y}^{2}/2)-x\theta_{x}-y\theta_{y}$. We consider the motion in a static magnetic field. According to conventional particle tracking the magnitude of the velocity is constant and is equal $v=ds/dt=\mathrm{const.}$, where $s(z)$ is the longitudinal coordinate along the path. The transverse components of the envelope of the field in Eq. (LABEL:rrevwied) in the far zone and paraxial approximation finally becomes
@@ 19.3 Undulator Radiation
Traditionally, all courses in synchrotron radiation theory have begun in the same way, retracing the path followed in the historical development of the subject. One first learns about synchrotron radiation from bending magnets. We will begin this chapter by dealing with the ”advanced” subject of the synchrotron radiation theory. For undulator setup covariant calculations of the radiation process are quite simple.
To generate specific synchrotron radiation characteristics, radiation is often produced from special insertion devices called undulators. The resonance approximation, which can always be applied in the case of undulator radiation setups, yields simplifications of the theory. This approximation does not replace the paraxial one, but it is used together with it. It takes advantage of another parameter that is usually large, i.e. number of undulator periods $N_{w}\gg 1$. The frequency emitted by a particle going through an undulator can be obtained by considering the interference between the parts of the radiation created at successive periods (see Fig. 54). The frequency of this field is Doppler shifted. The shortest wavelength is observed on the undulator axis. In resonance approximation, all undulator radiation at the shortest wavelength is emitted within an angle much smaller than $1/\gamma$ (see Fig. 55). This automatically selects observation angles of interest. If we consider observation angles outside the diffraction angle, we obtain zero intensity with accuracy $1/N_{w}$.
We now have to ask: Why did the error in insertion device theory remain so long undetected? We answer this question in great detail below. In discussing this we will deal only with the radiation into the central cone (see Fig. 55). In doing so, we neglected all information about the electron trajectory. In this approximation, the electron oscillation amplitude is always (independently of how large the undulator strength parameter is) much smaller than the radiation diffraction size at the undulator exit because $N_{w}\gg 1$. Thus, the undulator radiation theory gives fields very much like the (dipole-like) instantaneous theory. So, we are satisfied using the conventional (i.e. non-covariant) approach for describing the undulator radiation into the central cone, that is the practical situation of interest.
@@ 19.3.1 Conventional Theory
Eq. (43) can be used to characterize the far field from an electron moving on any path. In this section, we present a simple derivation of the frequency representation of the radiated field produced by an electron in the planar undulator. The magnetic field on the undulator axis has the form
@@
$m$ being the electron mass at rest and $B_{w}$ being the maximal magnetic field of the undulator on the axis.
In this case, the electron path is given by
@@
Usually, it does not make sense to calculate the intensity distribution from Eq. (51) alone, without extra-terms (both interfering and not) from the other parts of the electron path. This means that one should have complete information about the electron path and calculate extra terms to be added to Eq. (51) to have the total field from a given setup. Yet, we can find particular situations for which the contribution from Eq. (51) is dominant with respect to others. In this case Eq. (51), alone, has independent physical meaning.
@@
A well-known expression for the angular distribution of the first harmonic field in the far zone (see Appendix A2 for a detailed derivation) can be obtained from Eq. (51). Such expression is axis-symmetric, and can, therefore, be presented as a function of a single observation angle $\theta$, where $\theta^{2}=\theta_{x}^{2}+\theta_{y}^{2}$. One obtains the following distribution for the slowly varying envelope of the electric field:
@@
where $\mathrm{sinc}(\cdot)\equiv\sin(\cdot)/(\cdot)$. Therefore, the field is horizontally polarized and azimuthal symmetric. Eq. (60) describes a field with a spherical wavefront centered in the middle of the undulator.
@@ 19.3.2 Why did the Error in Insertion Device Theory Remain so Long Undetected?
We have seen that in all generality the expression for the undulator field in the far zone and in the ultrarelativistic (i.e. paraxial) approximation can be written as Eq. (141). Within the resonance approximation ($N_{w}\gg 1$) for the frequencies around the first harmonic, it can be simplified to the well-known expression Eq. (60) where the field is horizontally polarized and azimuthal symmetric. The divergence of this radiation is much smaller compared to the angle $1/\bar{\gamma}_{z}$. The mathematical reason stems from the fact that the factor $\sin(\cdot)/(\cdot)$ represents the well-known resonance character of the undulator radiation. If we are interested in the angular width of the peak around the observation angle $\theta=0$, we can introduce an angular displacement $\Delta\theta$. Taking the first zero of the $\sin(\cdot)/(\cdot)$ function at $C=0$ we will be able to determine the natural angular width of the radiation for the first harmonic $\theta_{c}$. The cone with aperture $\theta_{c}$ is usually called the central cone. It can be found that $\theta_{c}^{2}=1/(2N_{w}\bar{\gamma}_{z}^{2})\ll 1/\bar{\gamma}_{z}^{2}$.
@@
Now we would like to understand what is the characteristic transverse size of the field distribution at the exit of the undulator. The radiation from magnetic poles always interferes coherently at zero angle with respect to the undulator axis. This interference is constructive within an angle of about $\sqrt{c/(\omega L_{w})}$. We can estimate the interference size at the undulator exit as about $\sqrt{cL_{w}/\omega}$. On the other hand, the electron oscillating amplitude is given by $r_{w}=c\theta_{s}/k_{w}=cK/(\gamma k_{w})$. It follows that $r_{w}^{2}/(cL_{w}/\omega)=K^{2}\omega/(L_{w}k_{w}^{2}\gamma^{2})=K^{2}/[\pi N_{w}(1+K^{2}/2)]\ll 1$, where we use the fact that $\gamma^{2}=(1+K^{2}/2)\bar{\gamma}_{z}^{2}$. This inequality holds independently of the value of $K$ because $N_{w}\gg 1$. Thus, the electron oscillating amplitude is always much smaller than the radiation diffraction size at the undulator exit.
@@
We consider the radiation associated with the first order term in the expansion of the Eq. (147) in the power of small parameter $[K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq. (150)). But in doing so we miss all information about transverse electron trajectory in the phase factor Eq. (53) since the term $K\theta_{x}\omega\cos(k_{w}z^{\prime})/(\gamma k_{w}c)$ is neglected. In this approximation, the electron orbit scale is always much smaller than the radiation diffraction size, and Eq. (60) gives fields very much in agreement with the dipole radiation theory. So we are satisfied using the non covariant approach when considering the transverse electron motion.
@@
There are several points to be made about the above result. We have just explained that in accounting only for the radiation in the central cone, we miss all information about the transverse electron motion. To be complete we must add an analysis of the accelerated motion along the $z$-direction (i.e. along the undulator axis). We assume that the transverse velocity $v_{\perp}(z)$ is small compared to the velocity of light $c$. We consider the small expansion parameter $v_{\perp}/c$, neglecting terms of order $(v_{\perp}/c)^{3}$, but not of order $(v_{\perp}/c)^{2}$. In other words, we use a second-order relativistic approximation for transverse motion. We should remark that the analysis of the longitudinal motion in the ultrarelativistic approximation is much simpler than in the case of transverse motion. It is easy to see that the acceleration in the constant magnetic field yields a transverse electron velocity $v_{\perp}$ and $\Delta v_{z}=-v(v_{\perp}/c)^{2}/2$ parallel to the $z$-axis.
@@ 19.3.3 Influence of the Kick According to Conventional Theory
Eq. (60) can be generalized to the case of a particle with a given offset $\vec{l}$ and deflection angle $\vec{\eta}$ with respect to the longitudinal axis, assuming that the magnetic field in the undulator is independent of the transverse coordinate of the particle (see Appendix A3). Although this can be done using Eq. (161) directly, it is sometimes possible to save time by getting the answer with some trick. For example, in the undulator case, one takes advantage of the following geometrical considerations, which are in agreement with rigorous mathematical derivation. First, we consider the effect of an offset $\vec{l}$ on the transverse plane, with respect to the longitudinal axis $z$. Since the magnetic field experienced by the particle does not change, the far-zone field is simply shifted by a quantity $\vec{l}$. Eq. (60), can be immediately generalized by systematic substitution of the transverse coordinate of observation, $\vec{r}_{0}$ with $\vec{r}_{0}-\vec{l}$. This means that $\vec{\theta}=\vec{r}_{0}/z_{0}$ must be substituted by $\vec{\theta}-\vec{l}/z_{0}$, thus yielding
@@
Let us now discuss the effect of a deflection angle $\vec{\eta}$. Since the magnetic field experienced by the electron is assumed to be independent of its transverse coordinate, the path followed is still sinusoidal, but the effective undulator period is now given by $\lambda_{w}/\cos(\eta)\simeq(1+\eta^{2}/2)\lambda_{w}$. This induces a relative red shift in the resonant wavelength $\Delta\lambda/\lambda\sim\eta^{2}/2$. In practical cases of interest, we may estimate $\eta\sim 1/\gamma$. Then, $\Delta\lambda/\lambda\sim 1/\gamma^{2}$ should be compared with the relative bandwidth of the resonance, that is $\Delta\lambda/\lambda\sim 1/N_{w}$, $N_{w}$ being the number of undulator periods. For example, if $\gamma>10^{3}$, the red shift due to the deflection angle can be neglected in all situations of practical relevance. As a result, the introduction of a deflection angle only amounts to a rigid rotation of the entire system. Performing such rotation we should account for the fact that the phase factor in Eq. (61) is indicative of a spherical wavefront propagating outwards from position $z=0$ and remains thus invariant under rotations. The argument in the $\mathrm{sinc}(\cdot)$ function in Eq. (61), instead, is modified because the rotation maps the point $(z_{0},0,0)$ into the point $(z_{0},-\eta_{x}z_{0},-\eta_{y}z_{0})$. As a result, after rotation, Eq. (61) transforms to
@@
Finally, in the far-zone case, we can always work in the limit for $l/z_{0}\ll 1$, which allows one to neglect the term ${\vec{l}}/{z_{0}}$ in the argument of the $\mathrm{sinc}(\cdot)$ function, as well as the quadratic term in $\omega l^{2}/(2cz_{0})$ in the phase. Thus Eq. (63) can be further simplified, giving the generalization of Eq. (60) in its final form:
@@ 19.3.4 Influence of the Kick According to Correct Coupling of Fields and Particles
According to the correct coupling of fields and particles, there is a remarkable prediction of undulator radiation theory concerning the undulator radiation from the single electron with and without kick. Namely, when a kick is introduced, there is a red shift in the resonance wavelength of the undulator radiation in the velocity direction. To show this, let us consider the covariant treatment, which makes explicit use of Lorentz transformations.
@@
Suppose that without a kick the electron moves along the constrained trajectory parallel to the undulator axis. The field that produces this electron in the far zone is given by Eq. (60). Referring back to Eq. (64), we see that the conventional undulator radiation theory gives the following expression for radiation field after the kick
@@
This formula has nearly, but not quite the same form as Eq. (66), the difference consisting in the term $\omega\theta_{k}^{2}/(2c)$ in the argument of $\mathrm{sinc}$ function. Attention must be paid to the difference in resonance frequency between the undulator radiation setup with and without the kick. Remembering the definition of the detuning parameter $C=k_{w}\Delta\omega/\omega_{r}$, we can write the redshift in resonance frequency as $\Delta\omega/\omega_{r}=-\omega_{r}\theta_{k}^{2}/(2k_{w}c)$. With this, we also pointed out that the red shift can be written as $\Delta\omega/\omega_{r}=-\gamma^{2}\theta_{k}^{2}/(1+K^{2}/2)$. We now see a second order correction $\theta_{k}^{2}$ that is, however, multiplied by the large factor $\bar{\gamma}_{z}^{2}$.
We are now ready to investigate, more generally, what form the field expression takes under the introduction of a kick. Suppose that, without the kick, the electron moves along the trajectory with angle $\vec{\eta}$ with respect to the undulator axis. The field produced by this electron is given by Eq. (64). We let $\vec{\theta}_{k}$ be the kick angle of the electron with respect to its initial motion. The conventional approach gives the following expression for the field after the kick
@@
Now this all leads to an interesting situation. According to conventional theory, the resonance wavelength depends only on the observation angle with respect to the electron velocity direction. Equation (70) says that for any kick angle $\vec{\theta}_{k}$ and for any angle $\vec{\eta}$ between the undulator axis and the initial electron velocity direction, the radiation along the velocity direction has no redshift. We would like to emphasize a very important difference between conventional and covariant theory. The result of the covariant approach Eq. (72) depends on the absolute value of the kick angle $\theta_{k}$ and the radiation along the velocity direction has the redshift only when the kick angle has a nonzero value.
We must conclude that when we accelerate the electron in the lab frame upstream of the undulator, the information about this acceleration is included in the covariant trajectory.